OpenStudy (anonymous):

ra has 28¢ in his pocket. He has 20 coins that are either pennies or nickels. How many nickels does Ira have? Explain and show your work. *

OpenStudy (welshfella):

how many cents are there in a nickel and a penny.? (i'm from the UK)

OpenStudy (anonymous):

a penny is one cent and a nickle is 10 cents

OpenStudy (anonymous):

oops nicke is 5 cents

OpenStudy (welshfella):

ok so if there are x pennies and y nickels then we have the equation x + 10y = 28

OpenStudy (welshfella):

x + 5y = 28

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

now we have to fit while numbers to that equation

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

oh also he has 20 coins so that is x + y = 20 so we need to solve the system x + y = 20 x + 5y = 28

OpenStudy (welshfella):

now if we subtract equation 1 from equation 2 the x will disappear and w'ell get 4y = 8

OpenStudy (anonymous):

so first i would multiply then add right?

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

you have 4 times the number of nickels = 8 so divide by 4

OpenStudy (welshfella):